OpenStudy (anonymous):

Find the derivative of the limit process: f(x)=8/(x-3) Apparently I've forgotten how to do math soo I'm stuck evaluating this here: lim Δx->0 ((8/x+Δx-3) - (8/x-3))/Δx

OpenStudy (anonymous):

differentials calculus??

OpenStudy (amistre64):

\[f(x)=\frac{8}{x-3}; \ \ f(x+h)=\frac{8}{x+h-3}\] \begin{align} \lim_{h->0}&\frac{f(x+h)-f(x)}{h}\\\\ &\frac{\cfrac{8}{x+h-3}-\cfrac{8}{x-3}}{h}\\\\ &\frac{\cfrac{8(x-3)-8(x+h-3)}{(x+h-3)(x-3)}}{h}\\\\ &\frac{8(x-3)-8(x+h-3)}{(x+h-3)(x-3)(h)}\\\\ &\frac{8x-24-8x-8h+24}{(x+h-3)(x-3)(h)}\\\\ &\frac{-8h}{(x+h-3)(x-3)(h)}\\\\ &\frac{-8}{(x+h-3)(x-3)}\\\\ &\frac{-8}{(x+0-3)(x-3)}\\\\ &\frac{-8}{(x-3)^2}\\\\ \end{align}

OpenStudy (amistre64):

what did i miss?

myininaya (myininaya):

\[\lim_{\Delta x \rightarrow 0}\frac{\frac{8}{x+\Delta x-3}-\frac{8}{x-3}}{\Delta x} \]

myininaya (myininaya):

wait it didn't show all that just a sec ago openstudy is going so slow tonight

OpenStudy (amistre64):

8 -8*1 Dx ------ = -------- ... nothing l0ol (x-3) (x-3)^2

OpenStudy (amistre64):

im a ninja ;)

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

Thank you so much! Any website I can use to practice simple math. I'm so rusty and it will be my downfall on my first calc exam.

OpenStudy (anonymous):

you can practice here..

OpenStudy (anonymous):

And give ppl wrong answers ._.

OpenStudy (amistre64):

interactmath.com